# Latent Factor Analysis in Short Panels

### v2

@@ E.5.3 Feasible CLT for the FA estimators
Let us now establish a feasible CLT for $W_{n}$. In order to estimate matrix $Q_{\beta}$ in the asymptotic variance $\Omega_{W}$ in Lemma 7, we use the estimated betas and residuals, and combine them with a temporal sample splitting approach to cope with the EIV problem caused by the fixed $T$ setting. Specifically, let us split the time spell into two consecutive sub-intervals with $T_{1}$ and $T_{2}$ observations, with $T_{1}+T_{2}=T$ and such that $T_{1}>k$ and $T_{2}\geq k$. The factor model in the two sub-intervals reads $y_{1,i}=\mu_{1}+F_{1}\beta_{i}+\varepsilon_{1,i}$ and $y_{2,i}=\mu_{2}+F_{2}\beta_{i}+\varepsilon_{2,i}$, and let $V_{1,\varepsilon}$ and $V_{2,\varepsilon}$ denote the corresponding diagonal matrices of error average unconditional variances. [^41] We can take the two sub-intervals as the halves of the time span. If this choice does not meet conditions $T_{1}>k$ and $T_{2}\geq k$ in a subperiod, we take the second sub-interval such that $T_{2}=k$, and add to the first sub-interval a sufficient number of dates from the preceeding subperiod in order to get $T_{1}=k+1$. The conditions $T_{1}>k$ and $T_{2}\geq k$ are needed because we estimate residuals and betas in the first and the second sub-intervals, namely $\hat{\varepsilon}_{1,i}=M_{\hat{F}_{1},\hat{V}_{1,\varepsilon}}(y_{1,i}-\bar{y}_{1})$ and $\hat{\beta}_{i}=(\hat{F}_{2}^{\prime}\hat{V}_{2,\varepsilon}^{-1}\hat{F}_{2})^{-1}\hat{F}_{2}^{\prime}\hat{V}_{2,\varepsilon}^{-1}(y_{2,i}-\bar{y}_{2})$. Here, $\hat{F}_{j}$ and $\hat{V}_{j,\varepsilon}$ for $j=1,2$ are deduced from the FA estimates in the full period of $T$ observations. Define $\hat{\Psi}_{\beta}=\frac{1}{n}\sum_{m}\sum_{i,j\in I_{m}}(\hat{\beta}_{i}\hat{\beta}_{j}^{\prime})\otimes(\hat{\varepsilon}_{1,i}\hat{\varepsilon}_{1,j}^{\prime})$. By using $\hat{\varepsilon}_{1,i}=(M_{\hat{F}_{1},\hat{V}_{1,\varepsilon}}F_{1})\beta_{i}+M_{\hat{F}_{1},\hat{V}_{1,\varepsilon}}(\varepsilon_{1,i}-\bar{\varepsilon}_{1})$, $M_{\hat{F}_{1},\hat{V}_{1,\varepsilon}}F_{1}=O_{p}(\frac{1}{\sqrt{n}})$ and $\frac{1}{n^{2}}\sum_{m}b_{m,n}^{2}=\sum_{m}B_{m,n}^{2}=o(1)$, we get $\hat{\Psi}_{\beta}=(I_{k}\otimes M_{\hat{F}_{1},\hat{V}_{1,\varepsilon}})\left(\frac{1}{n}\sum_{m}\sum_{i,j\in I_{m}}(\hat{\beta}_{i}\hat{\beta}_{j}^{\prime})\otimes[(\varepsilon_{1,i}-\bar{\varepsilon}_{1})(\varepsilon_{1,j}-\bar{\varepsilon}_{1})^{\prime}]\right)(I_{k}\otimes M_{\hat{F}_{1},\hat{V}_{1,\varepsilon}}^{\prime})+o_{p}(1)$. Now, we use $\hat{\beta}_{i}=\left[(\hat{F}_{2}^{\prime}\hat{V}_{2,\varepsilon}^{-1}\hat{F}_{2})^{-1}\hat{F}_{2}^{\prime}\hat{V}_{2,\varepsilon}^{-1}F_{2}\right]\beta_{i}+(\hat{F}_{2}^{\prime}\hat{V}_{2,\varepsilon}^{-1}\hat{F}_{2})^{-1}\hat{F}_{2}^{\prime}\hat{V}_{2,\varepsilon}^{-1}(\varepsilon_{2,i}-\bar{\varepsilon}_{2})$, and $\bar{\varepsilon}_{1}=o_{p}(n^{-1/4})$, $\bar{\varepsilon}_{2}=o_{p}(n^{-1/4})$ from Lemma 5 (a), as well as the the mds condition in Assumption 3. We get $\hat{\Psi}_{\beta}=\hat{\Psi}_{\beta,1}+\hat{\Psi}_{\beta,2}+o_{p}(1)$, where $\hat{\Psi}_{\beta,1}=(I_{k}\otimes M_{F_{1},V_{1,\varepsilon}})\left(\frac{1}{n}\sum_{m}\sum_{i,j\in I_{m}}(\beta_{i}\beta_{j}^{\prime})\otimes(\varepsilon_{1,i}\varepsilon_{1,j}^{\prime})\right)(I_{k}\otimes M_{F_{1},V_{1,\varepsilon}}^{\prime})$ and $\hat{\Psi}_{\beta,2}=\left([(F_{2}^{\prime}V_{2,\varepsilon}^{-1}F_{2})^{-1}F_{2}^{\prime}V_{2,\varepsilon}^{-1}]\otimes M_{F_{1},V_{1,\varepsilon}}\right)\left(\frac{1}{n}\sum_{m}\sum_{i,j\in I_{m}}(\varepsilon_{2,i}\varepsilon_{2,j}^{\prime})\otimes(\varepsilon_{1,i}\varepsilon_{1,j}^{\prime})\right)\break\left([(F_{2}^{\prime}V_{2,\varepsilon}^{-1}F_{2})^{-1}F_{2}^{\prime}V_{2,\varepsilon}^{-1}]\right.\left.\otimes M_{F_{1},V_{1,\varepsilon}}\right)^{\prime}$. We use $\frac{1}{n}\sum_{m}\sum_{i,j\in I_{m}}(\beta_{i}\beta_{j}^{\prime})\otimes(\varepsilon_{1,i}\varepsilon_{1,j}^{\prime})=Q_{\beta}\otimes V_{1,\varepsilon}+o_{p}(1)$, and $\frac{1}{n}\sum_{m}\sum_{i,j\in I_{m}}(\varepsilon_{2,i}\varepsilon_{2,j}^{\prime})\otimes(\varepsilon_{1,i}\varepsilon_{1,j}^{\prime})=\Omega_{21}+o_{p}(1)$, where $\Omega_{21}$ is the sub-block of matrix $\Omega_{Z}$ that is the asymptotic variance of $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\varepsilon_{2,i}\otimes\varepsilon_{1,i}\Rightarrow N(0,\Omega_{21})$. Then, $\hat{\Psi}_{\beta}=Q_{\beta}\otimes(M_{F_{1},V_{1,\varepsilon}}V_{1,\varepsilon})+\left([(F_{2}^{\prime}V_{2,\varepsilon}^{-1}F_{2})^{-1}F_{2}^{\prime}V_{2,\varepsilon}^{-1}]\otimes M_{F_{1},V_{1,\varepsilon}}\right)\Omega_{21}\left([(F_{2}^{\prime}V_{2,\varepsilon}^{-1}F_{2})^{-1}F_{2}^{\prime}V_{2,\varepsilon}^{-1}]\otimes M_{F_{1},V_{1,\varepsilon}}\right)^{\prime}+o_{p}(1)$. Thus, we get a consistent estimator of $Q_{\beta}\otimes(V_{1,\varepsilon}^{-1/2}M_{F_{1},V_{1,\varepsilon}}V_{1,\varepsilon}^{1/2})$ by subtracting to $\hat{\Psi}_{\beta}$ a consistent estimator of the second term on the RHS (bias term), [^42] Sample splitting makes the estimation of the bias easier, but we can avoid such a splitting at the expense of a more complicated debiasing procedure. and then by pre- and post-multiplying times $(I_{k}\otimes\hat{V}_{1,\varepsilon}^{-1/2})$. To get a consistent estimator of $Q_{\beta}$, we apply a linear transformation that amounts to computing the trace of the second term of a Kronecker product, and divide by $Tr(V_{1,\varepsilon}^{-1/2}M_{F_{1},V_{1,\varepsilon}}V_{1,\varepsilon}^{1/2})=T_{1}-k$. Thus: $\hat{Q}_{\beta}=\frac{1}{n(T_{1}-k)}\sum_{m}\sum_{i,j\in I_{m}}(\hat{\beta}_{i}\hat{\beta}_{j}^{\prime})(\hat{\varepsilon}_{1,j}^{\prime}\hat{V}_{1,\varepsilon}^{-1}\hat{\varepsilon}_{1,i})-\frac{1}{T_{1}-k}\sum_{j=1}^{T_{1}}(I_{k}\otimes e_{j}^{\prime})\left\{\left([(\hat{F}_{2}^{\prime}\hat{V}_{2,\varepsilon}^{-1}\hat{F}_{2})^{-1}\hat{F}_{2}^{\prime}\hat{V}_{2,\varepsilon}^{-1}]\otimes\right.\right.\break\left.\left.[\hat{V}_{1,\varepsilon}^{-1/2}M_{\hat{F}_{1},\hat{V}_{1,\varepsilon}}]\right)\hat{\Omega}_{21}\left([\hat{V}_{2,\varepsilon}^{-1}\hat{F}_{2}(\hat{F}_{2}^{\prime}\hat{V}_{2,\varepsilon}^{-1}\hat{F}_{2})^{-1}]\otimes[M_{\hat{F}_{1},\hat{V}_{1,\varepsilon}}^{\prime}\hat{V}_{1,\varepsilon}^{-1/2}]\right)\right\}(I_{k}\otimes e_{j})$, where the $e_{j}$ are $T_{1}$-dimensional unit vectors, and $\hat{\Omega}_{21}$ is obtained from Subsection E.5.3 i). If estimate $\hat{Q}_{\beta}$ is not positive definite, we regularize it by deleting the negative eigenvalues.
@@
To get a feasible CLT for the FA estimators from (E.20)-(E.21), we need the joint distribution of the Gaussian matrix variates $Z$ and $W$. Under the condition of Lemma 7 (b), the estimates of the asymptotic variances of $vech(Z)$ and $vec(W)$ are enough, since these vectors are independent. Otherwise, to estimate the covariance $Cov(vech(Z),vec(W))$, we need to extend the approaches of the previous subsections.
E.5.4 Special cases
@@
Let us consider the case $\varepsilon_{i}\overset{ind}{\sim}(0,\sigma_{ii}V_{\varepsilon})$ where $V_{\varepsilon}=\bar{\sigma}^{2}I_{T}$, with independent components across time and the normalization $\underset{n\rightarrow\infty}{\lim}\frac{1}{n}\sum_{i}\sigma_{ii}=1$. By repeating the arguments of Section E.3 for the constrained FA estimators (see Section 4.3), we get $Tr(M_{F}(\Psi_{y}-\Psi_{\varepsilon})M_{F})=0$ instead of equation (E.4). It yields the asymptotic expansions $\sqrt{n}(\hat{\sigma}^{2}-\tilde{\sigma}^{2})=\frac{1}{T-k}Tr(M_{F}Z_{n})+o_{p}(1)=\frac{\bar{\sigma}^{2}}{T-k}Tr(Z_{n}^{*})+o_{p}(1)$, and $\sqrt{n}(\hat{F}_{j}-F_{j})=\frac{1}{\bar{\sigma}^{2}}R_{j}(\Psi_{y}-\Psi_{\varepsilon})F_{j}-\frac{1}{\bar{\sigma}^{2}}\Lambda_{j}\Psi_{\varepsilon}F_{j}+o_{p}(1)=\frac{1}{\bar{\sigma}^{2}}R_{j}(W_{n}F^{\prime}+FW_{n}^{\prime}+\bar{Z}_{n})F_{j}+o_{p}(1),$ where we use $\Psi_{y}-\Psi_{\varepsilon}=W_{n}F^{\prime}+FW_{n}^{\prime}+\bar{Z}_{n}$, $\Psi_{\varepsilon}=\frac{1}{T-k}Tr(M_{F}Z_{n})I_{T}$ and $\Lambda_{j}F_{j}=0$, and $\bar{Z}_{n}=Z_{n}-\frac{1}{T-k}Tr(M_{F}Z_{n})I_{T}$. Moreover, by sphericity, we have $R_{j}=\frac{1}{2\gamma_{j}}P_{F_{j}}+\frac{1}{\gamma_{j}}M_{F}+\sum_{\ell=1,\ell\neq j}^{k}\frac{1}{\gamma_{j}-\gamma_{\ell}}P_{F_{\ell}}$. Thus, we get $\sqrt{n}(\hat{\sigma}^{2}-\tilde{\sigma}^{2})\Rightarrow\frac{\bar{\sigma}^{2}}{T-k}Tr(Z^{*})$ and $\sqrt{n}(\hat{F}_{j}-F_{j})\Rightarrow\frac{1}{\bar{\sigma}^{2}}R_{j}(WF^{\prime}+FW^{\prime}+\bar{Z})F_{j}$. [^43] The asymptotic distribution of estimator $\hat{\sigma}^{2}$ coincides with that derived in FGS with perturbation theory methods. The asymptotic distribution of the factor estimates slightly differs from that given in FGS, Section 5.1, because of the different factor normalization adopted by FA compared to PCA even under sphericity. The Gaussian matrix $Z$ is such that $Z_{tt}\sim N(0,\eta)$ and $Z_{t,s}\sim N(0,q)$ for $t\neq s$, mutually independent, where $\eta=\underset{n\rightarrow\infty}{\lim}\frac{1}{n}\sum_{i}V[\varepsilon_{i,t}^{2}]$, and $vec(W)\sim N(0,Q_{\beta}\otimes I_{T})$. Variables $Z$ and $W$ are independent if $E[\varepsilon_{i,t}^{3}]=0$. FGS, Section 4.3.1, explain how we can estimate $q$ and $\eta$ by solving a system of two linear equations based on estimated moments of $\hat{\varepsilon}_{i,t}$.
@@
Let us now make the link with the distributional results in Anderson and Amemiya (1988). In our setting, the analogous conditions as those in their Corollary 2 would be: (a) random effects for the loadings that are i.i.d. with $E[\beta_{i}]=0$, $V[\beta_{i}]=I_{k}$, (b) error terms are i.i.d. $\varepsilon_{i}\sim(0,V_{\varepsilon})$ with $V_{\varepsilon}=diag(V_{\varepsilon,11},...,V_{\varepsilon,TT})$ such that $E[\varepsilon_{i,t}\varepsilon_{i,r}\varepsilon_{i,s}\varepsilon_{i,p}]=V_{\varepsilon,tt}V_{\varepsilon,ss}$, for $t=r>s=p$, and $=0$, otherwise, and (c) $\beta_{i}$ and $\varepsilon_{i}$ are mutually independent. Thus, $\sigma_{ii}=1$ for all $i$, i.e., errors are cross-sectionally homoskedastic. Under the aforementioned Conditions (a)-(c), the Gaussian distributional limits $Z$ and $W$ are such that $V[Z_{tt}]=\eta_{t}V_{\varepsilon,tt}^{2}$, for $\eta_{t}:=V[\varepsilon_{i,t}^{2}]/V_{\varepsilon,tt}^{2}$, $V[Z_{ts}]=V_{\varepsilon,tt}V_{\varepsilon,ss}$, for $t\neq s$, all covariances among different elements of $Z$ vanish, and $V[vec(W)]=I_{k}\otimes V_{\varepsilon}$. Equations (E.20)-(E.21) yield the asymptotic distributions of the FA estimates. In particular, they do not depend on the distribution of the $\beta_{i}$. Moreover, the distribution of the out-of-diagonal elements of $Z$ does not depend on the distribution of the errors, while, for the diagonal term, we have $\eta_{t}=2$ for Gaussian errors. As remarked in Section E.5.2, if the asymptotic distribution of estimator $\hat{V}_{\varepsilon}$ is centered around the realized matrix $\frac{1}{n}\sum_{i}\varepsilon_{i}\varepsilon_{i}^{\prime}$ instead of its expected value, that distribution involves the out-of-diagonal elements of $Z$, and the elements of $W$. Hence, in that case, the asymptotic distribution of the FA estimates is the same independent of the errors being Gaussian or not, and depends on $F$ and ${V}_{\varepsilon}$ only, as found in Anderson and Amemiya (1988).
E.6 Orthogonal transformations and maximal invariant statistic

### v3

@@ B.5.3 Feasible CLT for the FA estimators
Let us now establish a feasible CLT for $W_{n}$. In order to estimate matrix $Q_{\beta}$ in the asymptotic variance $\Omega_{W}$ in Lemma 8, we use the estimated betas and residuals, and combine them with a temporal sample splitting approach to cope with the EIV problem caused by the fixed $T$ setting. Specifically, let us split the time spell into two consecutive sub-intervals with $T_{1}$ and $T_{2}$ observations, with $T_{1}+T_{2}=T$ and such that $T_{1}>k$ and $T_{2}\geq k$. The factor model in the two sub-intervals reads $y_{1,i}=\mu_{1}+F_{1}\beta_{i}+\varepsilon_{1,i}$ and $y_{2,i}=\mu_{2}+F_{2}\beta_{i}+\varepsilon_{2,i}$, and let $V_{1,\varepsilon}$ and $V_{2,\varepsilon}$ denote the corresponding diagonal matrices of error average unconditional variances. [^14] We can take the two sub-intervals as the halves of the time span. If this choice does not meet conditions $T_{1}>k$ and $T_{2}\geq k$ in a subperiod, we take the second sub-interval such that $T_{2}=k$, and add to the first sub-interval a sufficient number of dates from the preceeding subperiod in order to get $T_{1}=k+1$. The conditions $T_{1}>k$ and $T_{2}\geq k$ are needed because we estimate residuals and betas in the first and the second sub-intervals, namely $\hat{\varepsilon}_{1,i}=M_{\hat{F}_{1},\hat{V}_{1,\varepsilon}}(y_{1,i}-\bar{y}_{1})$ and $\hat{\beta}_{i}=(\hat{F}_{2}^{\prime}\hat{V}_{2,\varepsilon}^{-1}\hat{F}_{2})^{-1}\hat{F}_{2}^{\prime}\hat{V}_{2,\varepsilon}^{-1}(y_{2,i}-\bar{y}_{2})$. Here, $\hat{F}_{j}$ and $\hat{V}_{j,\varepsilon}$ for $j=1,2$ are deduced from the FA estimates in the full period of $T$ observations. Define $\hat{\Psi}_{\beta}=\frac{1}{n}\sum_{m}\sum_{i,j\in I_{m}}(\hat{\beta}_{i}\hat{\beta}_{j}^{\prime})\otimes(\hat{\varepsilon}_{1,i}\hat{\varepsilon}_{1,j}^{\prime})$. By using $\hat{\varepsilon}_{1,i}=(M_{\hat{F}_{1},\hat{V}_{1,\varepsilon}}F_{1})\beta_{i}+M_{\hat{F}_{1},\hat{V}_{1,\varepsilon}}(\varepsilon_{1,i}-\bar{\varepsilon}_{1})$, $M_{\hat{F}_{1},\hat{V}_{1,\varepsilon}}F_{1}=O_{p}(\frac{1}{\sqrt{n}})$ and $\frac{1}{n^{2}}\sum_{m}b_{m,n}^{2}=\sum_{m}B_{m,n}^{2}=o(1)$, we get $\hat{\Psi}_{\beta}=(I_{k}\otimes M_{\hat{F}_{1},\hat{V}_{1,\varepsilon}})\left(\frac{1}{n}\sum_{m}\sum_{i,j\in I_{m}}(\hat{\beta}_{i}\hat{\beta}_{j}^{\prime})\otimes[(\varepsilon_{1,i}-\bar{\varepsilon}_{1})(\varepsilon_{1,j}-\bar{\varepsilon}_{1})^{\prime}]\right)(I_{k}\otimes M_{\hat{F}_{1},\hat{V}_{1,\varepsilon}}^{\prime})+o_{p}(1)$. Now, we use $\hat{\beta}_{i}=\left[(\hat{F}_{2}^{\prime}\hat{V}_{2,\varepsilon}^{-1}\hat{F}_{2})^{-1}\hat{F}_{2}^{\prime}\hat{V}_{2,\varepsilon}^{-1}F_{2}\right]\beta_{i}+(\hat{F}_{2}^{\prime}\hat{V}_{2,\varepsilon}^{-1}\hat{F}_{2})^{-1}\hat{F}_{2}^{\prime}\hat{V}_{2,\varepsilon}^{-1}(\varepsilon_{2,i}-\bar{\varepsilon}_{2})$, and $\bar{\varepsilon}_{1}=o_{p}(n^{-1/4})$, $\bar{\varepsilon}_{2}=o_{p}(n^{-1/4})$ from Lemma 6 (a), as well as the the mds condition in Assumption 11. We get $\hat{\Psi}_{\beta}=\hat{\Psi}_{\beta,1}+\hat{\Psi}_{\beta,2}+o_{p}(1)$, where $\hat{\Psi}_{\beta,1}=(I_{k}\otimes M_{F_{1},V_{1,\varepsilon}})\left(\frac{1}{n}\sum_{m}\sum_{i,j\in I_{m}}(\beta_{i}\beta_{j}^{\prime})\otimes(\varepsilon_{1,i}\varepsilon_{1,j}^{\prime})\right)(I_{k}\otimes M_{F_{1},V_{1,\varepsilon}}^{\prime})$ and $\hat{\Psi}_{\beta,2}=\left([(F_{2}^{\prime}V_{2,\varepsilon}^{-1}F_{2})^{-1}F_{2}^{\prime}V_{2,\varepsilon}^{-1}]\otimes M_{F_{1},V_{1,\varepsilon}}\right)\left(\frac{1}{n}\sum_{m}\sum_{i,j\in I_{m}}(\varepsilon_{2,i}\varepsilon_{2,j}^{\prime})\otimes(\varepsilon_{1,i}\varepsilon_{1,j}^{\prime})\right)\penalty-10000\left([(F_{2}^{\prime}V_{2,\varepsilon}^{-1}F_{2})^{-1}F_{2}^{\prime}V_{2,\varepsilon}^{-1}]\right.\left.\otimes M_{F_{1},V_{1,\varepsilon}}\right)^{\prime}$. We use $\frac{1}{n}\sum_{m}\sum_{i,j\in I_{m}}(\beta_{i}\beta_{j}^{\prime})\otimes(\varepsilon_{1,i}\varepsilon_{1,j}^{\prime})=Q_{\beta}\otimes V_{1,\varepsilon}+o_{p}(1)$, and $\frac{1}{n}\sum_{m}\sum_{i,j\in I_{m}}(\varepsilon_{2,i}\varepsilon_{2,j}^{\prime})\otimes(\varepsilon_{1,i}\varepsilon_{1,j}^{\prime})=\Omega_{21}+o_{p}(1)$, where $\Omega_{21}$ is the sub-block of matrix $\Omega_{Z}$ that is the asymptotic variance of $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\varepsilon_{2,i}\otimes\varepsilon_{1,i}\Rightarrow N(0,\Omega_{21})$. Then, $\hat{\Psi}_{\beta}=Q_{\beta}\otimes(M_{F_{1},V_{1,\varepsilon}}V_{1,\varepsilon})+\left([(F_{2}^{\prime}V_{2,\varepsilon}^{-1}F_{2})^{-1}F_{2}^{\prime}V_{2,\varepsilon}^{-1}]\otimes M_{F_{1},V_{1,\varepsilon}}\right)\Omega_{21}\left([(F_{2}^{\prime}V_{2,\varepsilon}^{-1}F_{2})^{-1}F_{2}^{\prime}V_{2,\varepsilon}^{-1}]\otimes M_{F_{1},V_{1,\varepsilon}}\right)^{\prime}+o_{p}(1)$. Thus, we get a consistent estimator of $Q_{\beta}\otimes(V_{1,\varepsilon}^{-1/2}M_{F_{1},V_{1,\varepsilon}}V_{1,\varepsilon}^{1/2})$ by subtracting to $\hat{\Psi}_{\beta}$ a consistent estimator of the second term on the RHS (bias term), [^15] Sample splitting makes the estimation of the bias easier, but we can avoid such a splitting at the expense of a more complicated debiasing procedure. and then by pre- and post-multiplying times $(I_{k}\otimes\hat{V}_{1,\varepsilon}^{-1/2})$. To get a consistent estimator of $Q_{\beta}$, we apply a linear transformation that amounts to computing the trace of the second term of a Kronecker product, and divide by $Tr(V_{1,\varepsilon}^{-1/2}M_{F_{1},V_{1,\varepsilon}}V_{1,\varepsilon}^{1/2})=T_{1}-k$. Thus: $\hat{Q}_{\beta}=\frac{1}{n(T_{1}-k)}\sum_{m}\sum_{i,j\in I_{m}}(\hat{\beta}_{i}\hat{\beta}_{j}^{\prime})(\hat{\varepsilon}_{1,j}^{\prime}\hat{V}_{1,\varepsilon}^{-1}\hat{\varepsilon}_{1,i})-\frac{1}{T_{1}-k}\sum_{j=1}^{T_{1}}(I_{k}\otimes e_{j}^{\prime})\left\{\left([(\hat{F}_{2}^{\prime}\hat{V}_{2,\varepsilon}^{-1}\hat{F}_{2})^{-1}\hat{F}_{2}^{\prime}\hat{V}_{2,\varepsilon}^{-1}]\otimes\right.\right.\penalty-10000\left.\left.[\hat{V}_{1,\varepsilon}^{-1/2}M_{\hat{F}_{1},\hat{V}_{1,\varepsilon}}]\right)\hat{\Omega}_{21}\left([\hat{V}_{2,\varepsilon}^{-1}\hat{F}_{2}(\hat{F}_{2}^{\prime}\hat{V}_{2,\varepsilon}^{-1}\hat{F}_{2})^{-1}]\otimes[M_{\hat{F}_{1},\hat{V}_{1,\varepsilon}}^{\prime}\hat{V}_{1,\varepsilon}^{-1/2}]\right)\right\}(I_{k}\otimes e_{j})$, where the $e_{j}$ are $T_{1}$-dimensional unit vectors, and $\hat{\Omega}_{21}$ is obtained from Subsection B.5.3 i). If estimate $\hat{Q}_{\beta}$ is not positive definite, we regularize it by deleting the negative eigenvalues.
@@
To get a feasible CLT for the FA estimators from (B.20)-(B.21), we need the joint distribution of the Gaussian matrix variates $Z$ and $W$. Under the condition of Lemma 8 (b), the estimates of the asymptotic variances of $vech(Z)$ and $vec(W)$ are enough, since these vectors are independent. Otherwise, to estimate the covariance $Cov(vech(Z),vec(W))$, we need to extend the approaches of the previous subsections.
B.5.4 Special cases
@@
Let us consider the case $\varepsilon_{i}\overset{ind}{\sim}(0,\sigma_{ii}V_{\varepsilon})$ where $V_{\varepsilon}=\bar{\sigma}^{2}I_{T}$, with independent components across time and the normalization $\underset{n\rightarrow\infty}{\lim}\frac{1}{n}\sum_{i}\sigma_{ii}=1$. By repeating the arguments of Section B.3 for the constrained FA estimators (see Section 5.3), we get $Tr(M_{F}(\Psi_{y}-\Psi_{\varepsilon})M_{F})=0$ instead of equation (B.4). It yields the asymptotic expansions $\sqrt{n}(\hat{\sigma}^{2}-\tilde{\sigma}^{2})=\frac{1}{T-k}Tr(M_{F}Z_{n})+o_{p}(1)=\frac{\bar{\sigma}^{2}}{T-k}Tr(Z_{n}^{*})+o_{p}(1)$, and $\sqrt{n}(\hat{F}_{j}-F_{j})=\frac{1}{\bar{\sigma}^{2}}R_{j}(\Psi_{y}-\Psi_{\varepsilon})F_{j}-\frac{1}{\bar{\sigma}^{2}}\Lambda_{j}\Psi_{\varepsilon}F_{j}+o_{p}(1)=\frac{1}{\bar{\sigma}^{2}}R_{j}(W_{n}F^{\prime}+FW_{n}^{\prime}+\bar{Z}_{n})F_{j}+o_{p}(1),$ where we use $\Psi_{y}-\Psi_{\varepsilon}=W_{n}F^{\prime}+FW_{n}^{\prime}+\bar{Z}_{n}$, $\Psi_{\varepsilon}=\frac{1}{T-k}Tr(M_{F}Z_{n})I_{T}$ and $\Lambda_{j}F_{j}=0$, and $\bar{Z}_{n}=Z_{n}-\frac{1}{T-k}Tr(M_{F}Z_{n})I_{T}$. Moreover, by sphericity, we have $R_{j}=\frac{1}{2\gamma_{j}}P_{F_{j}}+\frac{1}{\gamma_{j}}M_{F}+\sum_{\ell=1,\ell\neq j}^{k}\frac{1}{\gamma_{j}-\gamma_{\ell}}P_{F_{\ell}}$. Thus, we get $\sqrt{n}(\hat{\sigma}^{2}-\tilde{\sigma}^{2})\Rightarrow\frac{\bar{\sigma}^{2}}{T-k}Tr(Z^{*})$ and $\sqrt{n}(\hat{F}_{j}-F_{j})\Rightarrow\frac{1}{\bar{\sigma}^{2}}R_{j}(WF^{\prime}+FW^{\prime}+\bar{Z})F_{j}$. [^16] The asymptotic distribution of estimator $\hat{\sigma}^{2}$ coincides with that derived in FGS with perturbation theory methods. The asymptotic distribution of the factor estimates slightly differs from that given in FGS, Section 5.1, because of the different factor normalization adopted by FA compared to PCA even under sphericity. The Gaussian matrix $Z$ is such that $Z_{tt}\sim N(0,\eta)$ and $Z_{t,s}\sim N(0,q)$ for $t\neq s$, mutually independent, where $\eta=\underset{n\rightarrow\infty}{\lim}\frac{1}{n}\sum_{i}V[\varepsilon_{i,t}^{2}]$, and $vec(W)\sim N(0,Q_{\beta}\otimes I_{T})$. Variables $Z$ and $W$ are independent if $E[\varepsilon_{i,t}^{3}]=0$. FGS, Section 4.3.1, explain how we can estimate $q$ and $\eta$ by solving a system of two linear equations based on estimated moments of $\hat{\varepsilon}_{i,t}$.
@@
Let us now make the link with the distributional results in Anderson and Amemiya (1988). In our setting, the analogous conditions as those in their Corollary 2 would be: (a) random effects for the loadings that are i.i.d. with $E[\beta_{i}]=0$, $V[\beta_{i}]=I_{k}$, (b) error terms are i.i.d. $\varepsilon_{i}\sim(0,V_{\varepsilon})$ with $V_{\varepsilon}=diag(V_{\varepsilon,11},...,V_{\varepsilon,TT})$ such that $E[\varepsilon_{i,t}\varepsilon_{i,r}\varepsilon_{i,s}\varepsilon_{i,p}]=V_{\varepsilon,tt}V_{\varepsilon,ss}$, for $t=r>s=p$, and =0=0, otherwise, and (c) $\beta_{i}$ and $\varepsilon_{i}$ are mutually independent. Thus, $\sigma_{ii}=1$ for all $i$, i.e., errors are cross-sectionally homoskedastic. Under the aforementioned Conditions (a)-(c), the Gaussian distributional limits $Z$ and $W$ are such that $V[Z_{tt}]=\eta_{t}V_{\varepsilon,tt}^{2}$, for $\eta_{t}:=V[\varepsilon_{i,t}^{2}]/V_{\varepsilon,tt}^{2}$, $V[Z_{ts}]=V_{\varepsilon,tt}V_{\varepsilon,ss}$, for $t\neq s$, all covariances among different elements of $Z$ vanish, and $V[vec(W)]=I_{k}\otimes V_{\varepsilon}$. Equations (B.20)-(B.21) yield the asymptotic distributions of the FA estimates. In particular, they do not depend on the distribution of the $\beta_{i}$. Moreover, the distribution of the out-of-diagonal elements of $Z$ does not depend on the distribution of the errors, while, for the diagonal term, we have $\eta_{t}=2$ for Gaussian errors. As remarked in Section B.5.2, if the asymptotic distribution of estimator $\hat{V}_{\varepsilon}$ is centered around the realized matrix $\frac{1}{n}\sum_{i}\varepsilon_{i}\varepsilon_{i}^{\prime}$ instead of its expected value, that distribution involves the out-of-diagonal elements of $Z$, and the elements of $W$. Hence, in that case, the asymptotic distribution of the FA estimates is the same independent of the errors being Gaussian or not, and depends on $F$ and ${V}_{\varepsilon}$ only, as found in Anderson and Amemiya (1988).
B.6 Orthogonal transformations and maximal invariant statistic